A frequency distribution is shown below. Complete parts​ (a) and​ (b). The number of televisions per household in a small town

Televisions 0 1 2 3

Households 27 443 729 1410

​(a) Use the frequency distribution to construct a probability distribution.

x

​P(x)

0

1

2

3
Answer #1

solution:-

given that

total households = 27+443+729+1410 = 2609

the frequency distribution to construct a probability distribution

x P(x)

0 - 27/2609 = 0.010

1 - 443/2609 = 0.170

2 - 729/2609 = 0.279

3 -1410/2609 = 0.540